What is the solution? 5x+12=3x-6
Answer by checkley71(8403)

5X+12=3X-6
5X-3X=-6-12
2X=-18
X=-18/2
X=-9 ANSWER.
PROOF:
5*-9+12=3*-9-6
-45+12=-27-6
-33=-33